20 Lansing Corporation reported net income of $67 million for last year. Depreciation expense totaled $15...

20

Lansing Corporation reported net income of $67 million for last year. Depreciation expense totaled $15 million and capital expenditures came to $6 million. Free cash flow is expected to grow at a rate of 4.2% for the foreseeable future. Lansing faces a 40% tax rate and has a 0.36 debt to equity ratio with $250 million (market value) in debt outstanding. Lansing's equity beta is 1.92, the risk-free rate is currently 5% and the market risk premium is estimated to be 6.2%. What is the current total value of Lansing's equity (in millions)?
Answer #1

The calculation is shown below:

Beta of Asset = Beta of Equity/(1+Debt Equity Ratio*(1-Tax Rate))

Beta of Asset = 1.92/(1+0.36*(1-40%)) = 1.58

Required Return = Risk Free Rate + Beta of Asset*Market Risk Premium

Required Return = 5% + 1.58 * 6.2% = 14.8%

Firm Value = Free Cash Flow*(1+Growth Rate)/(Required Return - Growth Rate)

Firm Value = ($67 + $15 - $5)*(1+4.2%)/(14.80% - 4.2%) = $756.92

Value of equity = $756.92- $250

= $506.92